# SmileyLlama: Modifying Large Language Models for Directed Chemical Space Exploration

### v3

@@ 1 Introduction
Chemical Language Models (CLMs)1 trained on string representations such as Simplified Molecular-Input Line-Entry System (SMILES) 2 and SELF-referencIng Embedded Strings (SELFIES) 3 have emerged as a useful tool for de novo generation of molecules, best exemplified by molecules relevant to pharmaceutical applications and drug discovery4, 5, 6, 7, 8, 9. Nearly all CLMs for molecular generation have been trained from scratch on large quantities of data such as ChEMBL10 and ZINC11 and using different model architectures, including variational autoencoders12, recurrent neural networks (RNNs) long short term memory (LSTM) cells 13, 14, generative pre-trained transformers (GPTs)15, 16 and structured state space sequence (S4) models17, 18. In addition, CLMs have achieved advances in chemical generation tasks through further downstream optimization of molecules with additional training or different model frameworks.7, 8, 19, 20, 21 A recent such example is the iMiner algorithm that operates on SELFIES and uses deep reinforcement learning (RL) with real-time 3D docking to generate molecules with high potential for inhibiting protein targets.21
@@
Language models are statistical models of probability distributions of units of language, and can be adapted to generate meaningful text by sampling from these distributions.22 The most recent advancement of language models have resulted from the training of scaled-up transformers 23 on massive amounts of data, resulting in the creation of Large Language Models (LLMs) such as GPT4 and Llama.24, 25, 26. While LLMs were originally created for applications in natural language processing, in recent years there have also been efforts to adapt LLMs to chemistry. Some studies have accessed frontier LLMs for the purpose of assisting research in the form of virtual lab members, or even performing research autonomously.27, 28, 29 Other studies have focused on training LLM architectures on chemical data which can translate between natural and chemical languages30, 31 such as finding a molecule that is a close match to a ground truth molecule from which the description is assessed by fingerprint tanimoto similarity.32 Beyond chemical dictionary lookups, LLMs have also been used to perform mutation and crossover for an evolutionary algorithm to explore chemical space.33 There have also been studies of LLMs to modify SMILES strings to change the properties of the molecules that they represent.34, 35, 36, 37 Others have taken inspiration from LLMs to design CLMs with new capabilities, such as the ability to respond to prompts through a transformer-based architecture37, 38. However, to our knowledge, no CLM derived from a pre-trained general purpose LLM foundation model has reached the performance exhibited by modern CLMs that are trained from scratch with chemical data.
@@
Here, we demonstrate that an open-weight LLM, Meta-Llama-3.1-8B-Instruct (which we will refer to as “Llama” for the remainder of the paper) 26, can be converted into a CLM for generative tasks in drug discovery. Using supervised fine-tuning (SFT) and Direct Preference Optimization (DPO)39 of the pre-trained Llama with SMILES strings derived from ChEMBL, SmileyLlama generates drug-like molecules with desirable properties specified in a user-defined prompt with relevance to medicinal chemistry. SmileyLlama not only uses the ability of LLMs to have their behavior guided by prompts, but also their ability to serve as a foundation model which can be fine-tuned to specialize in various text-generation tasks for molecule generation that we show can match or exceed the performance of modern CLMs trained directly on chemical data. We further demonstrate that SmileyLlama greatly improves the reinforcement learning component of iMiner to more efficiently explore chemical space to create molecules optimized for 3D binding to target proteins, illustrated with the SARS-Cov-2 (SARS2) main protease (MPro).40 While our dataset and subsequent analyses are created with drug discovery in mind as a downstream application, this general procedure can be extended to other chemical applications such as chemical synthesis41.
@@ 2.1 Supervised Fine-Tuning and Direct Preference Optimization
In order to steer the outputs of the pre-trained Llama model 26 for drug molecule generation we first use supervised fine-tuning (SFT), in which the weights of Llama are further optimized on generic drug-like properties (Section 3.1). We constructed our fine-tuning dataset with the SMILES strings of approximately 2 million molecules from ChEMBL Dataset v33 10. For each molecule in our dataset, we randomly picked a number of properties of pharmaceutical interests to calculate using RDKit 42. These properties include ranges of hydrogen-bond (H-bond) donors and acceptors, molecular weight (MW), log of the partition coefficient (logP), number of rotatable bonds, the fraction of $sp^{3}$ carbons (i.e., the number of $sp^{3}$ carbons divided by the total number of carbon atoms in the molecule, termed $Fsp^{3}$), the presence/absence of macrocycles, the presence/absence of covalent warhead-related SMARTS patterns, the presence/absence of at least one undesirable SMARTS pattern, a SMILES string representation of a BRICS single-pass substructure and the specification of a chemical formula.
@@
For numerical properties, we choose ranges that are relevant for medicinal chemistry. In many medicinal applications, molecules must have suitable characteristics of drugs related not just to their chemical architecture but also to relevant biological phenomena (e.g., oral absorption, metabolism, distribution). For example, molecules invented in the early stages of drug design are typically smaller in size and lipophilicity than their follow-on analogs and optimized drugs, so a user can specify that a model obeys the Rule of Three—a common rule of thumb for early small molecule drug design—when generating molecules43. As another example, we choose the TPSA ranges based on those which tend to be orally bioavailable or able to pass through the placenta or blood-brain-barrier44. If a drug need not meet these criteria (e.g., an injectable drug for a peripheral target may not require as specific a TPSA range) then a user interfacing with SmielyLlama should be able to adjust the range criterion or eliminate it. Further information about the specifics of these properties and the ranges we choose to specify during training can be found in the Supplementary Information in Section 1.
@@
After calculating and picking these properties for each SMILES string, we construct a prompt containing values of these properties, with the “correct” completion being the SMILES string that these properties were calculated from. For example, we used a prompt with a system instruction of You love and excel at generating SMILES strings of drug-like molecules and a user instruction of the form Output a SMILES string for a drug like molecule with the following properties: if properties are specified, or Output a SMILES string for a drug like molecule: if no properties are specified. This trains our model to generate molecules that have properties specified (or not) in the prompt. We structure the prompts used for SFT so that, during inference, the properties of generated molecules can be specified in the prompt. This lets users avoid having to downselect the vast majority of generated molecules for having the correct characteristics—instead, users can simply prompt SmileyLlama to provide molecules with the characteristics they desire. See Figure 1 and Algorithm 1 in the Supporting Information for further elaboration of SFT training.
@@
Figure 1: A visualization of the SFT workflow for Smiley-Llama. Given the Llama-3.1-8B-Instruct model 26, we used prompt-response pairs consisting of calculated molecular properties and completed SMILES strings to fine-tune Llama on SMILES strings completions, yielding SmileyLlama. Crucially, we construct the prompt for each example using properties calculated from the correct response (a SMILES string from ChEMBLv33).
@@
We also use Direct Preference Optimization (DPO)39 which also updates the weights of Llama such that the language model is more likely to produce strings designated as "winners" and less likely to produce strings designated as "losers". We use DPO to further reinforce our model’s ability to robustly generate molecules with properties specified in the prompt for more specific task-oriented goals (Sections 3.2 and 3.3). Algorithmically we prompt our SFT model to generate molecules with a given property, and sample several SMILES strings and use RDKit42 to assess whether they have properties in line with what the prompt requested. We then pair up molecules which correctly followed the prompt with those that don’t as winners and losers, respectively, then use a single epoch of DPO to improve the model’s results. See Algorithm 2 in SI for pseudocode of this scoring and pairing procedure.
@@
Finally we consider DPO for generating unique and valid ligands which bind to a specific protein target, Sars2-Mpro using the iMiner framework, the details of which have been published elsewhere 21. During our optimization process, we prompt SmileyLlama to provide molecules with a “High SARS2PRO” score. After every 2000 distinct molecules are generated per iteration, we score each molecule and randomly pair it with another. The molecule with the higher score is designated as the "winner" and the other as the "loser", and we construct a dataset with eight such pairings for each molecule, and then optimize SmileyLlama using DPO with respect to this dataset. We use the iMiner scoring function of three times the docking score plus a custom drug-likeness function to score all molecules per iteration. Invalid molecules were assigned a score of -10. This process is described in Algorithm S3 in the Supplementary Information.
@@ 2.2 Additional Training Details
We performed both SFT and DPO on Llama using the Axolotl Package 26, 45. For both SFT and DPO, we use the Low-Rank Adaptation (LoRA) applied to the linear layers of the model and FlashAttention with an Adam optimizer, cross-entropy loss, and a cosine learning rate scheduler with a maximum learning rate of $2\times 10^{-5}$.46, 47, 48, 49 All prompts were formatted according to the Alpaca instruction format 50, 51. Additional parameters for our training are a LoRA rank of 32, a LoRA alpha of 16, a LoRA dropout of 5%, and 10 warmup steps. For SFT, we trained for 1 epoch using a batch size of 64 on a single 4xA40 node for approximately 1 day with a validation set $\sim$5the size of the original. We also note that we randomized the SMILES string representation of each molecule, and we tokenized all SMILES strings with the Llama3 tokenizer26 when interfacing with SmileyLlama.
@@ 3.1 Comparison of SmileyLlama from SFT with other CLMs
To test the generative ability of SmileyLlama compared with other existing CLMs across a range of architectures, we used the GuacaMol suite 52 to benchmark the validity, uniqueness, and novelty of the molecules as shown in Table 1. Additionally, KL divergence and Frechet Chemnet Distance 53 based on the GuacaMol definition (FCDGuac) are used to analyze the distributional shifts from the ChEMBL training data for drug-like molecules more specifically.
Table 1: GuacaMol benchmarks comparing Llama, SmileyLlama, and some common CLM architectures trained on ChEMBL. The model benchmarks include valid chemical molecules, uniqueness and novelty with respect to the training set, and distribution similarity evaluated using KL divergence and Frechet ChemNet distance based on the GaucaMol definition ${}_{Guac}=exp(-0.2*FCD)$. All benchmarks were performed with 10,000 samples at a temeperature $T=$1.0 and a maximum of 256 new tokens for Llama and SmileyLlama. Due to the proprietary nature of Llama’s training data and the SMILES contained therein, we cannot assess novelty, KL-Divergence or FCDGuac for Llama’s zero-shot performance.
@@
VGAE-MCTS54 1.000 1.000 1.000 0.659 0.009
LSTM18 0.983 0.999 0.848 0.993 0.901
GPT18 0.915 1.000 0.978 0.977 0.826
S4 18 0.971 0.997 0.961 0.994 0.853
Llama zero-shot 0.693 0.458 N/A N/A N/A
Llama twenty-shot 0.463 1.000 0.952 0.921 0.081
SmileyLlama 0.960 1.000 0.988 0.970 0.553
@@
We first analyze the ability of Llama to produce molecules, relying only on its pre-trained knowledge (zero-shot), or by providing it with one or more examples from the ChEMBL database in the formulated prompt (Table 1 and Supplementary Table S1). We find that even without SFT or examples provided in a prompt the LLM is able to produce some valid and unique SMILES strings, but is generally poor even with variations in hyperparameters such as temperature ($T$) compared to other state-of-the-art CLMs. In Table 1 it is seen that SFT significantly improves SmileyLlama’s ability to generate drug-like molecules as good as the CLMs, although FCDGuac is fairly low compared to CLMs.
That said, we still find that SmileyLlama generates very good agreement with ChEMBL quantities across a diverse property set as seen in Figure 2. The UMAP visualization in Figure 2a finds that while there are small relative differences in densities, SmileyLlama generates molecules in every well-represented region of the chemical space of ChEMBL. Likewise, the distribution of molecular properties of interest to medicinal chemistry are in very good agreement between the SmileyLlama-generated molecules and ChEMBL molecules as shown in Figure 2b. Hence there are no obvious chemical regions of ChEMBL that are inaccessible to SmileyLlama. Furthermore, it has recently come to our attention that certain undesirable molecular scaffolds might be oversampled in CLM generative tasks.21 Supplementary Table S1 shows that although small percentages of these unviable chemical structures are present in the ChEMBL training data itself, all CLMs do not oversample these undesirable molecular scaffolds. Finally, while training was conducted at $T=1.0$, exploration of temperature used at inference on the Guacamol benchmark (Supplementary Fig. 1) suggests that this temperature is adequate for all tests described in the Results.
@@
Figure 2: Distribution comparisons for different properties of the generated molecules from SmileyLlama (blue) with molecules from the training dataset from ChEMBL (gold). (a) UMAP visualization of a random selection of 10,000 ChEMBL molecules and 10,000 SmileyLlama-generated molecules, using 15 neighbors and a minimum distance of 0.1. (b) The molecular properties considered are fraction of $sp^{3}$ hybridized carbons and heteroatoms, number of heavy atoms, number of H-bond donors and acceptors, number of aliphatic and aromatic rings and the maximum ring size, number of rotatable bonds, quantitative estimate of drug-likelihood (QED) value55, MW, approximate log partition coefficient between octanol and water (ALOGP)56, polarizable surface area (PSA) and topological PSA 57, and the number of structural alerts 58. All benchmarks were at a temperature $T=$1.0 and a maximum of 256 new tokens.

### v4

@@ 1 Introduction
Chemical Language Models (CLMs)grisoni_chemical_2023 trained on string representations such as Simplified Molecular-Input Line-Entry System (SMILES) smiles and SELF-referencIng Embedded Strings (SELFIES) selfies have emerged as a useful tool for de novo generation of molecules, best exemplified by molecules relevant to pharmaceutical applications and drug discoverymol-gen-bg0, mol-gen-bg1, mol-gen-bg2, mol-gen-bg3, mol-gen-bg4, mol-gen-bg5. Nearly all CLMs for molecular generation have been trained from scratch on large quantities of data such as ChEMBLgaulton_ChEMBL_2012 and ZINCTingle2023 and using different model architectures, including variational autoencodersVAE2018, recurrent neural networks (RNNs) long short term memory (LSTM) cells lstm, mol-gen-rnn, generative pre-trained transformers (GPTs)gpt, mol-gen-gpt0 and structured state space sequence (S4) modelss4, Ozcelik2024s4. In addition, CLMs have achieved advances in chemical generation tasks through further downstream optimization of molecules with additional training or different model frameworks.mol-gen-bg3, mol-gen-bg4, mol-gen-gpt1, mol-opt1, iminer A recent such example is the iMiner algorithm that operates on SELFIES and uses deep reinforcement learning (RL) with real-time 3D docking to generate molecules with high potential for inhibiting protein targets.iminer
@@
Language models are statistical models of probability distributions of units of language, and can be adapted to generate meaningful text by sampling from these distributions.880083 The most recent advancement of language models have resulted from the training of scaled-up transformers vaswani2023attentionneed on massive amounts of data, resulting in the creation of Large Language Models (LLMs)radfordllms such as the proprietary GPT-4openai_gpt-4_2024 and the open-weight Llama.dubey_llama_2024. While LLMs were originally created for applications in natural language processing, in recent years there have also been efforts to adapt LLMs to chemistry. Some studies have accessed frontier LLMs for the purpose of assisting research in the form of virtual lab members, or even performing research autonomously.Zheng2023, boiko_autonomous_2023, m_bran_augmenting_2024 Other studies have focused on training LLM architectures on chemical data which can translate between natural and chemical languagesmolT5, pei2023BioT5 such as finding a molecule that is a close match to a ground truth molecule from which the description is assessed by fingerprint tanimoto similarity.yu2024llasmol Beyond chemical dictionary lookups, LLMs have also been used to perform mutation and crossover for an evolutionary algorithm to explore chemical space.molleo There have also been studies of LLMs to modify SMILES strings to change the properties of the molecules that they represent.guevorguian_small_2024, mol-llm-engine-1, drugllm, wu_leveraging_2024 Others have taken inspiration from LLMs to design CLMs with new capabilities, such as the ability to respond to prompts through a transformer-based architecturewu_leveraging_2024, mol-llm-engine-2. However, to our knowledge, no CLM derived from a pre-trained general purpose LLM foundation model has reached the performance exhibited by modern CLMs that are trained from scratch with chemical data.
@@
Here, we demonstrate that an open-weight LLM, Meta-Llama-3.1-8B-Instruct (which we will refer to as “Llama” for the remainder of the paper) dubey_llama_2024, can be converted into a CLM for generative tasks in drug discovery. The fact that Llama is open-weight offers several benefits as a foundation model. In particular, this allows users to train and share adapters, to perform training and inference without needing to store potentially valuable data on a remote server, to have complete control over the hyperparameters and algorithms used for fine-tuning, and to perform interpretability analyses on the model weights. Using supervised fine-tuning (SFT) and Direct Preference Optimization (DPO)rafailov_direct_2024 of the pre-trained Llama with SMILES strings derived from ChEMBL, SmileyLlama generates drug-like molecules with desirable properties specified in a user-defined prompt with relevance to medicinal chemistry. SmileyLlama not only uses the ability of LLMs to have their behavior guided by prompts, but also their ability to serve as a foundation model which can be fine-tuned to specialize in various text-generation tasks for molecule generation that we show can match or exceed the performance of modern CLMs trained directly on chemical data. We further demonstrate that SmileyLlama greatly improves the reinforcement learning component of iMiner to more efficiently explore chemical space to create molecules optimized for 3D binding to target proteins, illustrated with the SARS-Cov-2 (SARS2) main protease (MPro).mpro_review While our dataset and subsequent analyses are created with drug discovery in mind as a downstream application, this general procedure can be extended to other chemical applications such as chemical synthesisSun2025.
@@ 2.1 Supervised Fine-Tuning and Direct Preference Optimization
In order to steer the outputs of the pre-trained Llama model dubey_llama_2024 for drug molecule generation we first use supervised fine-tuning (SFT), in which the weights of Llama are further optimized on generic drug-like properties (Section 3.1). We constructed our fine-tuning dataset with the SMILES strings of approximately 2 million molecules from ChEMBL Dataset v33 gaulton_ChEMBL_2012. For each molecule in our dataset, we randomly picked a number of properties of pharmaceutical interests to calculate using RDKit Landrum2016RDKit2016_09_4. These properties include ranges of hydrogen-bond (H-bond) donors and acceptors, molecular weight (MW), log of the partition coefficient (logP), number of rotatable bonds, the fraction of $sp^{3}$ carbons (i.e., the number of $sp^{3}$ carbons divided by the total number of carbon atoms in the molecule, termed $Fsp^{3}$), the presence/absence of macrocycles, the presence/absence of covalent warhead-related SMARTS patterns, the presence/absence of at least one undesirable SMARTS pattern, a SMILES string representation of a BRICS single-pass substructure and the specification of a chemical formula.
@@
For numerical properties, we choose ranges that are relevant for medicinal chemistry. In many medicinal applications, molecules must have suitable characteristics of drugs related not just to their chemical architecture but also to relevant biological phenomena (e.g., oral absorption, metabolism, distribution). For example, molecules invented in the early stages of drug design are typically smaller in size and lipophilicity than their follow-on analogs and optimized drugs, so a user can specify that a model obeys the Rule of Three—a common rule of thumb for early small molecule drug design—when generating moleculesjhoti_rule_2013. As another example, we choose the TPSA ranges based on those which tend to be orally bioavailable or able to pass through the placenta or blood-brain-barrierveber_molecular_2002. If a drug need not meet these criteria (e.g., an injectable drug for a peripheral target may not require as specific a TPSA range) then a user interfacing with SmielyLlama should be able to adjust the range criterion or eliminate it. Further specifics of these properties and the ranges we choose to specify during training can be found in the Supplementary Information in Section 1.
@@
After calculating and picking these properties for each SMILES string, we construct a prompt containing values of these properties, with the “correct” completion being the SMILES string that these properties were calculated from. For example, we used a prompt with a system instruction of You love and excel at generating SMILES strings of drug-like molecules and a user instruction of the form Output a SMILES string for a drug like molecule with the following properties: if properties are specified, or Output a SMILES string for a drug like molecule: if no properties are specified. We choose to create prompts that assign SmileyLlama the role of an AI which excels at producing SMILES strings due to the track record of role-promptingchenUnleashingPotentialPrompt2025; we chose this prompt format due to its balance between motivation and brevity. Each property has a 50% chance of being calculated and specified in the prompt so that the model can operate equally well during inference, whether or not any properties are specified. This trains our model to generate molecules that have properties specified (or not) in the prompt. We structure the prompts used for SFT so that, during inference, the properties of generated molecules can be specified in the prompt. This lets users avoid having to downselect the vast majority of generated molecules for having the correct characteristics—instead, users can simply prompt SmileyLlama to provide molecules with the characteristics they desire. See Figure 1 and Algorithm 1 in the Supporting Information for further elaboration of SFT training.
@@
Figure 1: A visualization of the SFT workflow for Smiley-Llama. Given the Llama-3.1-8B-Instruct model dubey_llama_2024, we used prompt-response pairs consisting of calculated molecular properties and completed SMILES strings to fine-tune Llama on SMILES strings completions, yielding SmileyLlama. Crucially, we construct the prompt for each example using properties calculated from the correct response (a SMILES string from ChEMBLv33).
@@
We also use Direct Preference Optimization (DPO)rafailov_direct_2024 which also updates the weights of Llama such that the language model is more likely to produce strings designated as "winners" and less likely to produce strings designated as "losers". We use DPO to further reinforce our model’s ability to robustly generate molecules with properties specified in the prompt for more specific task-oriented goals (Sections 3.2 and 3.3). Algorithmically we prompt our SFT model to generate molecules with a given property, and sample several SMILES strings and use RDKitLandrum2016RDKit2016_09_4 to assess whether they have properties in line with what the prompt requested. We then pair up molecules which correctly followed the prompt with those that don’t as winners and losers, respectively, then use a single epoch of DPO to improve the model’s results. See Algorithm 2 in SI for pseudocode of this scoring and pairing procedure.
@@
Finally we consider DPO for generating unique and valid ligands which bind to a specific protein target, Sars2-Mpro using the iMiner framework, the details of which have been published elsewhere iminer. During our optimization process, we prompt SmileyLlama to provide molecules with a “High SARS2PRO” score. After every 2000 distinct molecules are generated per iteration, we score each molecule and randomly pair it with another. The molecule with the higher score is designated as the "winner" and the other as the "loser", and we construct a dataset with eight such pairings for each molecule and then optimize SmileyLlama using DPO with respect to this dataset. We empirically found 8 pairs to work well; more than 8 pairs tended to lead to a lack of diversity, whereas fewer than 8 pairs resulted in less sample efficient optimization. We use a scoring function of three times the docking score plus the iMiner drug-likeness functioniminer to score all molecules per iteration. Invalid molecules were assigned a score of -10. This process is described in Algorithm S3 in the Supplementary Information.
@@ 2.2 Additional Training Details
We performed both SFT and DPO on Llama using the Axolotl Package dubey_llama_2024, noauthor_axolotl-ai-cloudaxolotl_nodate. For both SFT and DPO, we use the Low-Rank Adaptation (LoRA) applied to the linear layers of the model and FlashAttention with an Adam optimizer, cross-entropy loss, and a cosine learning rate scheduler with a maximum learning rate of $2\times 10^{-4}$.hu_lora_2021, dao2022flashattention, dao2023flashattention2, kingma_adam_2017 All prompts were formatted according to the Alpaca instruction format alpaca, wang-etal-2023-self-instruct. Additional parameters for our training are a LoRA rank of 32, a LoRA alpha of 16, a LoRA dropout of 5%, and 10 warmup steps. We inherit these hyperparameters from standard practice with Axolotl. For example, LoRA hyperparameters identical to these were used in the Hermes 3 SFTHermes3TechnicalReport. For SFT, we trained for 1 epoch using a batch size of 64 on a single 4xA40 node for approximately 32 hours with a validation set $\sim$5% the size of the original, an amount which would cost approximately \$53 in October 2025 on Vast.ai. We also note that we randomized the SMILES string representation of each molecule, and we tokenized all SMILES strings with the Llama3 tokenizerdubey_llama_2024 when interfacing with SmileyLlama.
@@ 3.1 Comparison of SmileyLlama from SFT with other CLMs
To test the generative ability of SmileyLlama compared with other existing CLMs across a range of architectures, we used the GuacaMol suite brown2019guacamol to benchmark the validity, uniqueness, and novelty of the molecules as shown in Table 1. Additionally, KL divergence and Frechet Chemnet Distance preuer2018frechet based on the GuacaMol definition (FCDGuac) are used to analyze the distributional shifts from the ChEMBL training data for drug-like molecules more specifically. The Fréchet ChemNet Distance has been introduced as a metric for chemical and biological relevance for drug discoverypreuer2018frechet, and we note that FCDGuac=exp(-0.2*FCD) compresses the FCD distances themselves for the convenience of defining a 0 to 1 value like the other Guacamol metricsbrown2019guacamol. Hence the scores should actually be considered from a log perspective such that an FCD below 5 (i.e. FCDGuac =0.37) is in strong agreement with distributions of drug-like properties. Only when the FCDGuac score drops by one to two orders of magnitude are they considered to be poor FCD scores. We recommend in the future that others report the more straightforward FCD distances themselves to avoid this interpretative confusion.
[…]
@@
GraphMCTSIwata2023 1.000 1.000 0.994 0.522 0.015
VGAE-MCTSIwata2023 1.000 1.000 1.000 0.659 0.009
AAEIwata2023 0.822 1.000 0.998 0.886 0.529
LSTMOzcelik2024s4 0.983 0.999 0.848 0.993 0.901
GPTOzcelik2024s4 0.915 1.000 0.978 0.977 0.826
S4 Ozcelik2024s4 0.971 0.997 0.961 0.994 0.853
Llama zero-shot 0.688 0.457 0.635 0.736 0.002
Llama twenty-shot 0.465 0.999 0.949 0.913 0.079
SmileyLlama 0.958 1.000 0.987 0.967 0.686
@@
We first analyze the ability of Llama to produce molecules, relying only on its pre-trained knowledge (zero-shot), or by providing it with one or more examples from the ChEMBL database in the formulated prompt (Table 1 and Supplementary Table S1). We find that without SFT or examples provided in a prompt the LLM is unable to produce high percentages of valid SMILES strings compared to other state-of-the-art CLMs, and is generally poor even with variations in hyperparameters such as temperature ($T$). Interestingly, validity is lower when 20 examples are provided in the prompt (twenty-shot) than it is when no examples are in the prompt (zero-shot). We speculate that Llama zero-shot has had some exposure to the SMILES syntax to be able to generate valid strings, but it has no intrinsic ability to generalize and it falls back to repeating the memorized SMILES and resulting in low uniqueness. When several examples are given in the prompt, this biases Llama away from the known SMILES strings it can produce, but the few examples means that its grasp on the allowed mutable structure of SMILES strings is poor and thus less valid. However, because these prompts are so diverse, Llama’s 20-shot uniqueness is very high.
In Table 1 it is seen that SFT significantly improves SmileyLlama’s ability to generate drug-like molecules as well as the CLMs. In addition, we experiment with the format of the SmileyLlama prompt, performing SFT on Llama with a less anthropomorphic user prompt and a blank template as an ablation study. We find that changing this prompt format does not significantly affect the guacamol benchmarks. More information on this prompt format, as well as that for the Llama models, can be found in Supplementary Information section 1.3 and Table S1. To show the generality of the LLM-SFT approach, we also fine-tune Llama-3.2-3B, Llama-3.2-1B, and Qwen-2.5-7Bqwen25TechnicalReport2025 using the same SFT workflow (including identical hyperparameters) that we developed for SmileyLlama, yielding the models we name SmileyLlama-1B, SmileyLlama-3B, and SmileyQwen2.5-7B, respectively. Supplementary Table S1 finds that the Guacamol benchmark results did not change substantially between SmileyLlama and SmileyQwen2.5-7B. We also find through inspecting the results of SmileyLlama-1B and SmileyLlama-3B, that the validity increases with parameter count, though novelty, uniqueness and the match between the training distribution and the distribution of generated molecules stays roughly unchanged. While our later analyses are restricted to SmileyLlama, these results suggest that a more capable, future LLMs could also be converted into a CLM using an identical or similar SFT method.
@@
Figure 2: Distribution comparisons for different properties of the generated molecules from SmileyLlama (blue) with molecules from the training dataset from ChEMBL (gold). (a) UMAP visualization of a random selection of 10,000 ChEMBL molecules and 10,000 SmileyLlama-generated molecules, using 15 neighbors and a minimum distance of 0.1; these are normal values in chemical space visualizationorlovHighDimensionsHuman2025. (b) The molecular properties considered are fraction of $sp^{3}$ hybridized carbons and heteroatoms, number of heavy atoms, number of H-bond donors and acceptors, number of aliphatic and aromatic rings and the maximum ring size, number of rotatable bonds, quantitative estimate of drug-likelihood (QED) valuebickerton2012quantifying, MW, approximate log partition coefficient between octanol and water (ALOGP)wildman1999prediction, polarizable surface area (PSA) and topological PSA Prasanna2009, and the number of structural alerts brenk2008lessons. All benchmarks were at a temperature T=T=1.0 and a maximum of 256 new tokens.
Figure 2 shows that SmileyLlama generates very good agreement with ChEMBL quantities across a diverse property set. The UMAP visualization in Figure 2a, a popular visualization tool used in drug discovery, finds that while there are small relative differences in densities, SmileyLlama generates molecules in every well-represented region of the chemical space of ChEMBL. Like all data reduction techniques, UMAP representations are qualitative, and hence we also consider the distribution of molecular properties of interest to medicinal chemistry in Figure 2b, and KL-divergence values indicates all properties are in very good agreement between the SmileyLlama-generated molecules and ChEMBL molecules. Hence there are no obvious chemical regions of ChEMBL that are inaccessible to SmileyLlama. Furthermore, it has recently come to our attention that certain undesirable molecular scaffolds might be oversampled in CLM generative tasks.iminer Supplementary Table S1 shows that although small percentages of these unviable chemical structures are present in the ChEMBL training data itself, all CLMs do not oversample these undesirable molecular scaffolds. Finally, while training was conducted at $T=1.0$, exploration of temperature used at inference on the Guacamol benchmark (Supplementary Fig. 1) suggests that this temperature is adequate for all tests described in the Results.